A particle of mass 2.5kg is suspended from a fixed point by a light inextensible string of length 1.8m.the particle moves in a horizontal circle on a smooth floor. calculate the angular speed of the particle if the tension in the string is 28 N

1
Expert's answer
2021-12-16T11:30:55-0500

Given:

"m=2.5\:\rm kg"

"l=1.8\:\rm m"

"F=28\:\rm N"


The Newton's second law says

"ma=F"

Using definition of the centripetal acceleration, we obtain

"m\frac{v^2}{l}=F"

Finally

"v=\sqrt{Fl/m}=\sqrt{28*1.8/2.5}=4.5\:\rm m/s"